# An Adaptive Method for Contextual Stochastic Multi-armed Bandits with Rewards Generated by a Linear Dynamical System

### v2

@@ Abstract
Index Terms:
Online learning, sequential decision-making, stochastic dynamical systems, stochastic multi-armed bandits, Kalman filters
@@ I Introduction
For this work, we consider the rewards generated by a Linear Gaussian Dynamical System (LGDS), and we envision it can be utilized to address challenges in many applications, including dynamic pricing problem [25, 26], controller selection for time-varying stochastic linear dynamical system [27, 28], opportunistic spectrum access in wireless networks [29, 15], and hyperparameter optimization [30] or algorithm selection [31] for machine learning models. Each of the applications above can be modeled as a dynamical system as previous efforts have made a similar assumption such as [26] for dynamic pricing, [32] for hyperparameter optimization, and [27] for controller selection. Finally, many of the examples above have a context available, i.e. the computed loss from a cost function in a machine learning model can be viewed as a context.
@@
Notation: For any $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$, we have the inner product $\left\langle x,y\right\rangle=x^{\top}y\in\mathbb{R}$. For the natural number $i\in\mathbb{N}$, we have $[i]=\{1,2,\dots,i\}$.
@@ II Problem Statement
$\begin{cases}z_{t+1}&=\Gamma z_{t}+\xi_{t},~{}z_{0}\sim\mathscr{N}(0,\Sigma_{0})\\ \theta_{t}&=C_{\theta}z_{t}+\phi_{t}\\ X_{t}&=\langle c_{a},z_{t}\rangle+\eta_{t}\end{cases},$ (1)
where $z_{t}\in\mathbb{R}^{d}$ is the state of the system and $\Gamma\in\mathbb{R}^{d\times d}$ is the system state matrix. For each round $t=1,2,\dots,n$ and $n>d$, the learner observes the reward $X_{t}\in\mathbb{R}$ based on the chosen action and the context $\theta_{t}\in\mathbb{R}^{m}$. The context $\theta_{t}$ is a dynamically changing value with respect to the LGDS state variable $z_{t}$ that the learner always observes and its observation matrix $C_{\theta}\in\mathbb{R}^{m\times d}$ is constant. The noises $\xi_{t}\in\mathbb{R}^{d}$, $\phi_{t}\in\mathbb{R}^{m}$, and $\eta_{t}\in\mathbb{R}$ are i.i.d. normally distributed, i.e., $\xi_{t}\sim\mathscr{N}(\mu,Q)$ with $Q\succeq 0$, $\phi_{t}\sim\mathscr{N}(0,R)$ with $R\succ 0$, and $\eta_{t}\sim\mathscr{N}(0,\sigma_{\eta}^{2})$ with $\sigma_{\eta}>0$. For the noise term $\eta_{t}\in\mathbb{R}$, we will assume that each action $c_{a}\in\mathcal{A}$ has its own independently generated noise $\eta_{t}\in\mathbb{R}$.
@@ III Matrix Representation of the System
where $\hat{z}_{t|t}\triangleq\mathbb{E}\left[z_{t}|\mathscr{F}_{t}\right]$, $P_{t|t}\triangleq\mathbb{E}\left[(z_{t}-\hat{z}_{t|t})(z_{t}-\hat{z}_{t|t})^{\top}|\mathscr{F}_{t}\right]$, and $\mathscr{F}_{t}$ is the sigma algebra generated by previous contexts $\theta_{0},\dots,\theta_{t}$. Since the pair $(\Gamma,C_{\theta})$ is observable (see Assumption 2), the Kalman gain matrix $K_{t}$ converges in few steps [33]. Therefore, the following steady-state Kalman filter can be used:
@@
Since system parameters are unknown, taking the inspiration from [33], we develop a method to learn the matrix representation of the Kalman filter. This matrix will be a function of the matrices and vectors defined in the steady-state Kalman filter given in (5) and (9). Let $s>0$ be the horizon length of how far we look into the past. From (5), we can express the reward prediction $\hat{X}_{t}$ using the observations from $t-s$ to $t-1$, as follows:
@@
According to (15), at round $t$, given the set of contexts from round $t-s$ up to round $t-1$ (i.e., $\{\theta_{t-s},\dots,\theta_{t-1}\}$), one can predict the reward $X_{t}$, where the prediction accuracy is impacted by the terms $\beta_{a}^{t}$ and $\varepsilon_{a}^{t}$ given in (16) and (11), respectively.
@@
According to discussions in [21], it is reasonable to assume that there exists a known constant $B_{G}$ such that $\|G_{a}\left(s\right)\|_{2}\leq B_{G}$ for all $a\in[k]$, $s\in\{1,\dots,s_{N}\}$.
@@
There exists a known constant $B_{c}>0$ such that $\left\|c_{a}\right\|_{2}\leq B_{c},~{}a\in\{1,\cdots,k\}$.
@@ IV-A Perturbation Method
$\bullet$ STEP 1: Providing a bound on the model error $\left\|\hat{G}_{a}^{t}\left(s\right)-G_{a}\left(s\right)\right\|_{V_{a}^{t}\left(s\right)}$.
$\bullet$ STEP 2: Providing a bound on the error $\left|\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}-G_{a}\left(s\right)^{\top}\Theta_{t}\right|$.
IV-A1 STEP 1
To provide a bound on the model error $\left\|\hat{G}_{a}^{t}\left(s\right)-G_{a}\left(s\right)\right\|_{V_{a}^{t}\left(s\right)}$, first we need to provide a bound on $\left\|\mathbf{B}_{\mathscr{T}_{a}}\right\|_{2}$. The lemma below will be used for proving the bound of $\left\|\mathbf{B}_{\mathscr{T}_{a}}\right\|_{2}$.
Lemma 1.
Let $X\in\mathbb{R}^{k}$ be a normally distributed random variable, i.e. $X\sim\mathcal{N}\left(0,\Sigma\right)$. The random variable $\sqrt{X^{\top}X}$ is bounded by a $\sqrt{\mbox{tr}\left(\Sigma\right)}$-subgaussian random variable.
Proof.
First, we will prove that $\sqrt{X^{\top}X}$ is bounded by a half-normally distributed random variable. Note the following equations
$\sqrt{X^{\top}X}$ $=\sqrt{\left(\Sigma^{1/2}Y\right)^{\top}\left(\Sigma^{1/2}Y\right)}$
$=\sqrt{Y^{\top}\Sigma Y}$
$=\sqrt{Y^{\top}V\Lambda V^{\top}Y}$
$=\sqrt{\left(V^{\top}Y\right)^{\top}\Lambda\left(V^{\top}Y\right)},$
where $Y\sim\mathcal{N}\left(0,I_{k}\right)$ and $\Lambda\triangleq\mbox{diag}\left(\lambda_{1},\dots,\lambda_{k}\right)$ is the diagonal matrix with the eigenvalues $\lambda_{1},\dots,\lambda_{k}$ associated with the eigenvectors $V=\left(v_{1},\dots,v_{k}\right)$ of the covariance matrix $\Sigma$. Since $\Sigma\succeq 0$, then the matrix $V$ is normal, i.e. $V^{\top}V=VV^{\top}=I_{k}$. Therefore, $Z\triangleq V^{\top}Y$, $Z\sim\mathcal{N}\left(0,I_{k}\right)$. This implies that
$\sqrt{X^{\top}X}=\sqrt{Z^{\top}\Lambda Z}=\sqrt{\sum_{i=1}^{k}\lambda_{i}Z_{i}^{2}}\overset{(a)}{\leq}\sum_{i=1}^{k}\sqrt{\lambda_{i}}\left|Z_{i}\right|.$
where in $(a)$ we used the triangle inequality. Therefore, $\sqrt{X^{\top}X}$ is bounded by a sum of half-normal distributed random variables $\left|Z_{i}\right|$. Since the half-normal random variable is subgaussian [35] and the sum of independent subgaussian random variables is subgaussian [1], we proved that $\sqrt{X^{\top}X}$ is bounded by a subgaussian random variable. To bound the variance of $\sqrt{X^{\top}X}$, we can just derive the variance of $\sum_{i=1}^{k}\lambda_{i}\left|Z_{i}\right|$:
$\mbox{Var}\left(\sum_{i=1}^{k}\sqrt{\lambda_{i}}\left|Z_{i}\right|\right)$ =$\overset{(b)}{=}\sum_{i=1}^{k}\lambda_{i}\mbox{Var}\left(\left|Z_{i}\right|\right)$
$=\mbox{tr}\left(\Sigma\right)\left(1-\frac{2}{\pi}\right),$
$\Rightarrow\mbox{Var}\left(\sum_{i=1}^{k}\sqrt{\lambda_{i}}\left|Z_{i}\right|\right)\leq\mbox{tr}\left(\Sigma\right),$
where in $(b)$ we used the fact that the variance of a sum of independent subgaussian random variables is the sum of variances [1]. Therefore, $\sqrt{X^{\top}X}$ is bounded by a $\sqrt{\mbox{tr}\left(\Sigma\right)}$-subgaussian, which completes the proof. ∎
Lemma 2.
$\left\|\mathbf{O}_{\mathscr{T}_{a}}\mathbf{B}_{\mathscr{T}_{a}}^{\top}\right\|_{V_{a}^{t}\left(s\right)^{-1}}\leq\\ \sqrt{2N_{a}\log\left(N_{a}/\delta\right)}B_{\beta}\left(s\right)\sqrt{\mbox{tr}\left(I_{ms+1}-\lambda V_{a}^{t}\left(s\right)^{-1}\right)},$ (25)
where $B_{\beta}\left(s\right)$ is defined to be
$B_{\beta}\left(s\right)\triangleq B_{c}\sqrt{\mbox{tr}\left(Z_{t-s}\right)}\left\|\left(\Gamma-\Gamma KC_{\theta}\right)^{s}\right\|_{2},$ (26)
and $B_{c}$ is as in Assumption 4.
Proof.
According to (22), we have:
$=$ $\left\|\sum_{t\in\mathscr{T}_{a}}V_{a}^{t}\left(s\right)^{-1/2}\Theta_{t}\beta_{a}^{t}\right\|_{2},$ (27)
$\left\|\mathbf{O}_{\mathscr{T}_{a}}\mathbf{B}_{\mathscr{T}_{a}}^{\top}\right\|_{V_{a}^{t}\left(s\right)^{-1}}\leq\sum_{t\in\mathscr{T}_{a}}\left|\beta_{a}^{t}\right|\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)^{-1}\Theta_{t}}.$ (28)
$\left|\beta_{a}^{t}\right|\leq\left\|c_{a}\right\|_{2}\left\|(\Gamma-\Gamma KC_{\theta})^{s}\right\|_{2}\left\|\hat{z}_{t-s}\right\|_{2},$ (29)
$\left|\beta_{a}^{t}\right|\leq B_{c}\left\|(\Gamma-\Gamma KC_{\theta})^{s}\right\|_{2}\sqrt{2\mathbb{E}\left[\left\|\hat{z}_{t-s}\right\|_{2}^{2}\right]\log\left(1/\delta\right)},$ (30)
$\left|\beta_{a}^{t}\right|\leq B_{\beta}\left(s\right)\sqrt{2\log\left(1/\delta\right)}.$ (31)
where $B_{\beta}\left(s\right)$ is defined in (26).
Regarding the term $\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)^{-1}\Theta_{t}}$ in (28), we have:
$~{}~{}\leq\sum_{t\in\mathscr{T}_{a}}B_{\beta}\left(s\right)\sqrt{2\log\left(1/\delta\right)}\sqrt{\mbox{tr}\left(V_{a}^{t}\left(s\right)^{-1}\Theta_{t}\Theta_{t}^{\top}\right)\cdot 1}$
$~{}~{}\overset{(b)}{\leq}B_{\beta}\left(s\right)\sqrt{2\log\left(1/\delta\right)}\sqrt{\sum_{t\in\mathscr{T}_{a}}\mbox{tr}\left(V_{a}^{t}\left(s\right)^{-1}\Theta_{t}\Theta_{t}^{\top}\right)}\cdot\sqrt{\sum_{t\in\mathscr{T}_{a}}1}$
$~{}~{}\leq\sqrt{N_{a}}B_{\beta}\left(s\right)\sqrt{2\log\left(1/\delta\right)}\sqrt{\sum_{t\in\mathscr{T}_{a}}\mbox{tr}\left(V_{a}^{t}\left(s\right)^{-1}\Theta_{t}\Theta_{t}^{\top}\right)}$
$~{}~{}\leq\sqrt{N_{a}}B_{\beta}\left(s\right)\sqrt{2\log\left(1/\delta\right)}\sqrt{\mbox{tr}\left(V_{a}^{t}\left(s\right)^{-1}\sum_{t\in\mathscr{T}_{a}}\Theta_{t}\Theta_{t}^{\top}\right)},$ (32)
where $(b)$ is concluded based on the Cauchy-Schwarz inequality. Since $\sum_{\tau\in\mathscr{T}_{a}}\Theta_{\tau}\Theta_{\tau}^{\top}=V_{a}^{t}\left(s\right)-\lambda I_{ms+1}$, we can rewrite (IV-A1) as follows which is satisfied with a probability of at least $1-N_{a}\delta$:
$\left\|\mathbf{O}_{\mathscr{T}_{a}}\mathbf{B}_{\mathscr{T}_{a}}^{\top}\right\|_{V_{a}^{t}\left(s\right)^{-1}}\leq\sqrt{2N_{a}\log\left(1/\delta\right)}B_{\beta}\left(s\right)\cdot\\ \sqrt{\mbox{tr}\left(V_{a}^{t}\left(s\right)^{-1}\left(V_{a}^{t}\left(s\right)-\lambda I_{ms+1}\right)\right)}.$ (33)
Note that the matrix $\sum_{\tau\in\mathscr{T}_{a}}\Theta_{\tau}\Theta_{\tau}^{\top}\succeq 0$, implying that $V_{a}^{t}\left(s\right)\succeq\lambda I_{ms+1}$ Thus, the terms in the square root of inequality (33) is always non-negative. According to (31) and (33) and setting $\delta=\delta/N_{a}$ no further effort is needed to show that inequality (25) is satisfied with a probability of at least $1-\delta$. ∎
Using Lemma 2, the following theorem completes STEP 1.
@@ IV-A1 STEP 1
$b_{a}^{t}\left(\delta,s\right)\triangleq$ $\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\det(V_{a}^{t}\left(s\right))^{1/2}}{\det(\lambda I)^{1/2}}\right)}$
$+\sqrt{2N_{a}\log\left(N_{a}/\delta\right)}B_{\beta}\left(s\right)\sqrt{\mbox{tr}\left(I-\lambda V_{a}^{t}\left(s\right)^{-1}\right)}$
$+\lambda B_{G}\sqrt{\mbox{tr}\left(V_{a}^{t}\left(s\right)^{-1}\right)}.$ (34)
Proof.
From (17) and (23), we have:
$+\left(\mathbf{B}_{\mathscr{T}_{a}}+\mathbf{E}_{\mathscr{T}_{a}}\right)\mathbf{O}_{\mathscr{T}_{a}}^{\top}V_{a}^{t}\left(s\right)^{-1}.$ (35)
Adding and subtracting the term $\lambda G_{a}\left(s\right)^{\top}V_{a}^{t}\left(s\right)^{-1}$ from the right-hand side of (35) yields:
$-\lambda G_{a}\left(s\right)^{\top}V_{a}^{t}\left(s\right)^{-1},$
$\hat{G}_{a}^{t}\left(s\right)^{\top}-G_{a}\left(s\right)^{\top}=$ $-\lambda G_{a}\left(s\right)^{\top}V_{a}^{t}\left(s\right)^{-1}$
$+\left(\mathbf{B}_{\mathscr{T}_{a}}+\mathbf{E}_{\mathscr{T}_{a}}\right)\mathbf{O}_{\mathscr{T}_{a}}^{\top}V_{a}^{t}\left(s\right)^{-1}.$ (36)
$\Big{(}\hat{G}_{a}^{t}\left(s\right)-G_{a}\left(s\right)\Big{)}^{\top}V_{a}^{t}\left(s\right)\\ =\left(-\lambda G_{a}\left(s\right)+\mathbf{O}_{\mathscr{T}_{a}}\left(\mathbf{B}_{\mathscr{T}_{a}}+\mathbf{E}_{\mathscr{T}_{a}}\right)^{\top}\right)^{\top}.$ (37)
Multiplying the right- and left-hand sides of (37) by transpose of the right- and left-hand sides of (37), respectively, and according to the fact that $V_{a}^{t}\left(s\right)^{-1}=\left(V_{a}^{t}\left(s\right)^{-1}\right)^{\top}$, we obtain the following:
$\Big{\|}\hat{G}_{a}^{t}\left(s\right)-G_{a}\left(s\right)\Big{\|}_{V_{a}^{t}\left(s\right)}\\ =\left\|-\lambda G_{a}\left(s\right)+\mathbf{O}_{\mathscr{T}_{a}}\left(\mathbf{B}_{\mathscr{T}_{a}}+\mathbf{E}_{\mathscr{T}_{a}}\right)^{\top}\right\|_{V_{a}^{t}\left(s\right)^{-1}}.$ (38)
$\left\|\hat{G}_{a}^{t}\left(s\right)-G_{a}\left(s\right)\right\|_{V_{a}^{t}\left(s\right)}\leq\left\|\lambda G_{a}\left(s\right)\right\|_{V_{a}^{t}\left(s\right)^{-1}}\\ +\left\|\mathbf{O}_{\mathscr{T}_{a}}\mathbf{B}_{\mathscr{T}_{a}}^{\top}\right\|_{V_{a}^{t}\left(s\right)^{-1}}+\left\|\mathbf{O}_{\mathscr{T}_{a}}\mathbf{E}_{\mathscr{T}_{a}}^{\top}\right\|_{V_{a}^{t}\left(s\right)^{-1}}.$ (39)
$\left\|\mathbf{O}_{\mathscr{T}_{a}}\mathbf{E}_{\mathscr{T}_{a}}^{\top}\right\|_{V_{a}^{t}\left(s\right)^{-1}}\leq\\ \sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\det(V_{a}^{t}\left(s\right))^{1/2}}{\det(\lambda I)^{1/2}}\right)}.$ (40)
Finally, according to Assumption 3, we have:
$\leq\lambda B_{G}\sqrt{\mbox{tr}\left(V_{a}^{t}\left(s\right)^{-1}\right)}.$ (41)
Thus, from Lemma 2, and according to inequalities (40) and (IV-A1), one can obtain the bound given in (1), which completes the proof. ∎
IV-A2 STEP 2
Using results from STEP 1, the following lemma will complete STEP 2.
@@ IV-A2 STEP 2
Proof.
Consider the following optimization problem
$\begin{array}[]{cc}\underset{\tilde{G}\in\mathbb{R}^{(ms+1)\times 1}}{\max}&\tilde{G}^{\top}\Theta_{t}\\ \mbox{ s.t. }&\tilde{G}\in\mathscr{C}_{a,s}\end{array},$ (43)
where the set $\mathscr{C}_{a,s}$ is defined as follows:
$\mathscr{C}_{a,s}=\left\{\tilde{G}\big{|}\left\|\hat{G}_{a}^{t}\left(s\right)-\tilde{G}\right\|_{V_{a}^{t}\left(s\right)}\leq b_{a}^{t}(\delta,s)\right\},$ (44)
with $b_{a}^{t}\left(\delta,s\right)$ as in (1). The Lagrangian $L(\tilde{G},\epsilon)$ for optimization problem (43) is
$L\left(\tilde{G},\epsilon\right)=\tilde{G}^{\top}\Theta_{t}\\ -\epsilon\left(\left\|\hat{G}_{a}^{t}\left(s\right)-\tilde{G}\right\|_{V_{a}^{t}\left(s\right)}^{2}-b_{a}^{t}\left(\delta,s\right)^{2}\right),$ (45)
where $\epsilon$ is Lagrange multiplier. The partial derivatives of $L\left(\tilde{G},\epsilon\right)$ are
$\frac{\partial L(\cdot)}{\partial\tilde{G}^{\top}}=\Theta_{t}^{\top}-2\epsilon\tilde{G}^{\top}V_{a}^{t}\left(s\right)+2\epsilon\hat{G}_{a}^{t}\left(s\right)^{\top}V_{a}^{t}\left(s\right),$
and
$\frac{\partial L(\cdot)}{\partial\epsilon}=-\left\|\hat{G}_{a}^{t}\left(s\right)-\tilde{G}\right\|_{V_{a}^{t}\left(s\right)}^{2}+b_{a}^{t}\left(\delta,s\right)^{2}.$ (46)
$\Theta_{t}^{\top}-2\epsilon\tilde{G}^{\top}V_{a}^{t}\left(s\right)+2\epsilon\hat{G}_{a}^{t}\left(s\right)^{\top}V_{a}^{t}\left(s\right)=0,$ (47)
which implies that
$\tilde{G}^{\top}=\hat{G}_{a}^{t}\left(s\right)^{\top}+\frac{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)^{-1}}{2\epsilon}.$ (48)
Using $\tilde{G}$ given in (48) in (46), and solving $\partial L\left(\tilde{G},\epsilon\right)/\partial\epsilon=0$ for $\epsilon$ yields:
$\epsilon=\frac{\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)^{-1}\Theta_{t}}}{2b_{a}^{t}\left(\delta,s\right)}.$ (49)
Plugging the obtained $\epsilon$ in (48) yields
$\tilde{G}^{\top}$ $=\hat{G}_{a}^{t}\left(s\right)^{\top}+\frac{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)^{-1}}{2\epsilon}$
which is the solution of the optimization problem (43). Since this optimal value lies on the edge of the confidence interval $\mathscr{C}_{a,s}$ given in (44), then with a probability of at least $1-2\delta$, we have
$G_{a}\left(s\right)^{\top}\Theta_{t}\leq\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}+b_{a}^{t}\left(\delta,s\right)\sqrt{\Theta_{t}^{\top}V_{a}^{t}\left(s\right)^{-1}\Theta_{t}},$
which yields the inequality (42) and completes the proof. ∎
IV-A3 STEP 3
Finally, since STEP 1 and STEP 2 have been proven, we can now design the perturbation signal $u_{a}$. A principle commonly used for bandits is Optimism in the Face of Uncertainty [1]. This principle states that the learner should be optimistic of actions it is uncertain of. The amount of optimism a learner has for each action is based on the error of the action’s reward prediction. Therefore, we will initially design $u_{a}$ based on this approach. For analysis, Theorem 2 and Corollary 1 investigate the impact of the perturbation signal $u_{a}$ on selecting the optimal action.

### v3

@@ I Introduction
For this work, we consider the rewards generated by a Linear Gaussian Dynamical System (LGDS), and we envision it can be utilized to address challenges in many applications, including dynamic pricing problem [25, 26], controller selection for time-varying stochastic linear dynamical system [27, 28], opportunistic spectrum access in wireless networks [29, 15], portfolio optimization [30, 31, 32], and hyperparameter optimization [33, 34] or algorithm selection [35] for machine learning models. Each of the applications above can be modeled as a dynamical system as previous efforts have made a similar assumption such as [26] for dynamic pricing, [36, 34] for hyperparameter optimization, and [27] for controller selection. Finally, many of the examples above have a context available, i.e. the computed loss from a cost function in a machine learning model can be viewed as a context.
@@
Notation: For any $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$, we have the inner product $\left\langle x,y\right\rangle=x^{\top}y\in\mathbb{R}$. For the natural number $i\in\mathbb{N}$, we have $i]=\{1,2,\dots,i\}$. The distribution $\mathscr{N}\left(\mu,\Sigma\right)$ is a normal distribution with a mean $\mu\in\mathbb{R}^{d}$ and a covariance of $\Sigma\succeq\mathbf{0}\in\mathbb{R}^{d\times d}$. The norm $\left\|x\right\|_{Q}\triangleq\sqrt{x^{\top}Qx}$ is the weighted $\ell$-2 norm where $Q\succ\mathbf{0}$.
@@ II Problem Statement
$\begin{cases}z_{t+1}&=\Gamma z_{t}+\xi_{t},~z_{0}\sim\mathscr{N}(\mathbf{0},\Sigma_{0})\\ \theta_{t}&=C_{\theta}z_{t}+\phi_{t}\\ X_{t}&=\langle c_{a},z_{t}\rangle+\eta_{t}\end{cases},$ (1)
where $z_{t}\in\mathbb{R}^{d}$ is the state of the system and $\Gamma\in\mathbb{R}^{d\times d}$ is the system state matrix. For each round $t=1,2,\dots,n$ and $n>d$, the learner observes the reward $X_{t}\in\mathbb{R}$ based on the chosen action and the context $\theta_{t}\in\mathbb{R}^{m}$. The context $\theta_{t}$ is a dynamically changing value with respect to the LGDS state variable $z_{t}$ that the learner always observes and its observation matrix $C_{\theta}\in\mathbb{R}^{m\times d}$ is constant. The noises $\xi_{t}\in\mathbb{R}^{d}$, $\phi_{t}\in\mathbb{R}^{m}$, and $\eta_{t}\in\mathbb{R}$ are i.i.d. normally distributed, i.e., $\xi_{t}\sim\mathscr{N}(\mu,Q)$ with $Q\succeq\mathbf{0}$, $\phi_{t}\sim\mathscr{N}(\mathbf{0},R)$ with $R\succ\mathbf{0}$, and $\eta_{t}\sim\mathscr{N}(0,\sigma_{\eta}^{2})$ with $\sigma_{\eta}>0$. For the noise term $\eta_{t}\in\mathbb{R}$, we will assume that each action $c_{a}\in\mathcal{A}$ has its own independently generated noise $\eta_{t}\in\mathbb{R}$.
@@ III Matrix Representation of the System
where
$\hat{z}_{t|t}$ $\triangleq\mathbb{E}\left[z_{t}|\mathscr{F}_{t}\right]$
$P_{t|t}$ $\triangleq\mathbb{E}\left[(z_{t}-\hat{z}_{t|t})(z_{t}-\hat{z}_{t|t})^{\top}|\mathscr{F}_{t}\right],$
and $\mathscr{F}_{t}$ is the sigma algebra generated by previous contexts $\theta_{0},\dots,\theta_{t}$. Since the pair $(\Gamma,C_{\theta})$ is observable (see Assumption 2), the Kalman gain matrix $K_{t}$ converges in few steps [37]. Therefore, the following steady-state Kalman filter can be used:
@@
Since system parameters are unknown, taking the inspiration from [37], we develop a method to learn the matrix representation of the Kalman filter. This matrix will be a function of the matrices and vectors defined in the steady-state Kalman filter given in (5) and (9). Let $s>0$ be the horizon length of how far we look into the past. From (5), we can express the reward prediction $\hat{X}_{t}$ using the observations from tst-s to t1t-1, as follows:
@@
According to (15), at round $t$, given the set of contexts from round tst-s up to round t1t-1 (i.e., $\{\theta_{t-s},\dots,\theta_{t-1}\}$), one can predict the reward $X_{t}$, where the prediction accuracy is impacted by the terms $\beta_{a}^{t}$ and $\varepsilon_{a}^{t}$ given in (16) and ($1$), respectively.
@@
According to discussions in [21], it is reasonable to assume that there exists a known constant $B_{G}$ such that $\|G_{a}\left(s\right)\|_{2}\leq B_{G}$ for all $a\in[k]$, $s\in[s_{N}]$.
@@
There exists a known constant $B_{c}>0$ such that $\left\|c_{a}\right\|_{2}\leq B_{c},~a\in[k]$.
@@ IV-A Perturbation Method
$\bullet$ STEP 1: Providing a bound on the model error
$\left\|\hat{G}_{a}^{t}\left(s\right)-G_{a}\left(s\right)\right\|_{V_{a}^{t}\left(s\right)}.$
This step provides Theorem 1 which is written below. All proofs can be found in the appendix of our Arxiv version [39].
@@
$b_{a}^{t}\left(\delta,s\right)\triangleq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\det(V_{a}^{t}\left(s\right))^{1/2}}{\det(\lambda I)^{1/2}}\right)}\\ +\sqrt{N_{a}g_{\Sigma_{z_{t}}}\left(\delta/N_{a}\right)}B_{\beta}\left(s\right)\sqrt{\mbox{tr}\left(I-\lambda V_{a}^{t}\left(s\right)^{-1}\right)}\\ +\lambda B_{G}\sqrt{\mbox{tr}\left(V_{a}^{t}\left(s\right)^{-1}\right)}.$ (25)
where $g_{\Sigma_{z_{t}}}\left(\delta\right)$ is defined to be
$g_{\Sigma_{z_{t}}}\left(\delta\right)\triangleq\mbox{tr}\left(\Sigma_{z_{t}}\right)+\\ \max\left\{\sqrt{\frac{K^{4}\left\|\Sigma_{z_{t}}\right\|_{HS}^{2}}{\tilde{c}}\log\left(\frac{2}{\delta}\right)},\frac{K^{2}\left\|\Sigma_{z_{t}}\right\|_{2}}{\tilde{c}}\log\left(\frac{2}{\delta}\right)\right\},$ (26)
$\tilde{c}>0$ is an absolute constant and $K=\sqrt{\left(1/2\right)\log\left(2\right)}$. The term $B_{\beta}\left(s\right)$ is defined to be
$B_{\beta}\left(s\right)\triangleq B_{c}\left\|\left(\Gamma-\Gamma KC_{\theta}\right)^{s}\right\|_{2}.$ (27)
The derivations and discussions for this bound are found in Appendix B.1.
$\bullet$ STEP 2: Providing a bound on the error
$\left|\hat{G}_{a}^{t}\left(s\right)^{\top}\Theta_{t}-G_{a}\left(s\right)^{\top}\Theta_{t}\right|.$
The bound for the prediction error above is Lemma 1 which is written below.
@@
The derivations and discussions for this bound are found in Appendix B.2.
IV-A1 STEP 3
Since STEP 1 and STEP 2 have been proven in Appendices B.1 and B.2, we can now design the perturbation signal $u_{a}$. A principle commonly used for bandits is Optimism in the Face of Uncertainty [1]. This principle states that the learner should be optimistic of actions it is uncertain of. The amount of optimism a learner has for each action is based on the error of the action’s reward prediction. Therefore, we will initially design $u_{a}$ based on this approach. For analysis, Theorem 2 and Corollary 1 investigate the impact of the perturbation signal $u_{a}$ on selecting the optimal action. Proofs for Theorem 2 and Corollary 1 are found in Appendix C. Extended versions of Theorem 2 and Corollary 1 are provided in Appendix C.